The perimeter is 32 inches it's length is 4 inches what is the width
A) 3 inches
B) 14
C) 28
D) 6

Answer by srinivas.g(540)

perimeter = 2 times the sum of length and width
= 2*( length+width)
32 = 2*(4 +width)
divide with 2 on both sides
32/2 = 2/2 * (4+width)
16 = 4 +width
subtract 4 on both sides
16-4 = 4+width -4
12 = width
width = 12 inches